how would you find out if the equation is parallel with the given line and it contains the given point
y=1/2x+5 (-6,1)

Answer by stanbon(75887)

how would you find out if the equation is parallel with the given line and it contains the given point
y=1/2x+5 (-6,1)
---------------------
Two lines are parallel if they have equal slopes
and unequal y-intercepts.
------------
(-6,1) is on y = (1/2)x+5 if
1 = (1/2)(-6)+5
1 = -3+5
1 = 2
----
That is not true so (-6,1) is not on the line y = (1/2)x+5
-----------------
Cheers,
Stan H.
-----------------